# The Anderson transition – a view from Krylov space

### v3

@@ I Introduction
Modeling quenched disorder has been instrumental in understanding the electronic properties of solids. In the case of the non-interacting Anderson model, spatial disorder is known to localize the electronic states if it is larger than a critical disorder strength, whereas below this strength the electron states are extended throughout the system [Anderson, Mott1968-3d, Kramer1993-localization_review, Evers_2008_AndersonTransitions]. Applying these concepts to interacting systems has been a major endeavor in the past decades [BASKO20061126, gornyi2005MBL, Oganesyan2007, annurev-conmatphys-031214-014726, Huse2014Phenomenoloy, Imbrie2016, ROS2015420, Serbyn2013, Thomson2018, AbaninReview, annurev-conmatphys-031214-014726, ALET2018498, Schreiber2015, Rubio-Abadal2019, Smith2016, Kohlert2019, Lukin2019, Rispoli2019, Pal2010MBLTransision, Lev2015, Luitz2015, Wahl2017, Mierzejewski2016, Mace2019, Chanda2020, Laflorencie2020, Gray2018, Doggen2018, POLFED2020, Serbyn2015MBLTransition, Sierant2020ThoulessTime, Chanda2020], but the interacting problem, dubbed many-body localization (MBL), has proved significantly more challenging both conceptually and numerically [suntajs2019quantum, suntajs2020transition, Dynamicalobstruction, sierant2025mbl].
@@
We may interpret these localized ergodicity-breaking systems to be effectively integrable, meaning they must possess an extensive set of local integrals of motion (LIOMs). Their existence may be motivated intuitively, invoking adiabatic evolution of the Hamiltonian from completely decoupled degrees of freedom [Imbrie2017_LIOMs], e.g. qubits, in which case LIOMs are Pauli-like operators (i.e. having discrete spectra $\pm 1$), or they can be obtained by simply time-averaging the Hamiltonian evolution of initially localized qubits [chandran2015constructing, geraedts2017emergent]. Though not uniquely defined, LIOMs should be exponentially localized with a small enough localization length to be stable against so called “quantum avalanches” induced by rare Griffith’s regions [Chandran2016_MBLbeyondEigenstates, deRoeck17, Luitz2017_HowaSmallQuantumBathCanThermalize, Ponte2017_HowOneSpin, MorningstarAvalanches, Bath-induceddelocalization, Embedded_Thermal_inclusion, Selsthermalizationofdiluteimpurities, Crowley2020_AvalancheInducedLocalizedandThermalRegions, Potirniche2019_ExplorationofStabilityofMBL, Rubio-Abadal2019, Crowley2022_MeanFieldTheoryofFailedThermalizingAvalanches, Suntajs2022_ErgodicityBreakinginZeroDimensions, Long2023_Prethermal_Phenom, Leonard2023avalanchesexperiment]. Given their non-unique construction and the stringent conditions on their decay, it is important to continue looking for new methods to understand the emergence of local integrals of motion. The Krylov subspace expansion, commonly known as the recursion method [recursionmethod_vis], is one such method which, although not a new technique, has been increasingly explored as a probe of quantum chaos (for a recent review, see Ref. [nandy2024quantumdynamicskrylovspace]). The LIOMs we will explore in this work are the same operators as the time-averaged LIOMs of Ref. [chandran2015constructing], but, importantly, expressed in the dynamically generated Krylov basis, which is arguably more natural and demonstrably more compact than any apriori fixed basis.
@@
The edge mode nature of LIOMs studied below has been identified in recent work on nearly integrable spin chains [Yates2020_LifetimeofAlmostStrong, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode]. By way of surveying other recent efforts to exploit Krylov methods, we note there have been many studies of quantum chaos and integrability in Krylov space [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Suchsland2025_KrylovComplexityandTrotterTransitions, Camargo2024_SpectralandKrylovComplexityBilliards, Cao_2021StatisticalmethodforOG, menzler2024krylovlocalizationprobeergodicity, baggioli2024_krylovcomplexityorderparameter, Yates2020DynamicsofAlmostStrongEdgeModes, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode, Yeh2024UniversalModelofFloquet, loizeau2025openingkrylovspaceaccess, Trigueros2022_KrylovComplexityofMBL, Balasubramanian2025_ChaosIntegrabilityandLateTimes, Yates2020_LifetimeofAlmostStrong, Scialchi2025_ExploringQuantumErgodicity_KrylovApproach, Uskov2024_QuantumDynamicsinOneandTwoDims, Teretenkov_2025_PseudomodeExpansion, Loizeau_2025_PauliStrings, Yeh2025_MomentMethodandContinuedFractionExpansion, Bartsch2024, angelinos2025temperaturedependencekrylovspace] which come with various proposals for detecting ergodicity breaking transitions. For one, chaos has been associated with a maximum growth rate of the Lanczos coefficients [Parker2019_universaloperatorgrowth], though this is not always the case [Bhattacharjee2022_KComplexity_SaddleDominatedScrambling]. The interplay between integrability, integrability breaking, and dimerization has been pointed out in Refs. [Rabinovici2022Krylovlocalizationandsuppression, Hashimoto2023_KrylovComplexityandChaos, nandy2024quantumdynamicskrylovspace, menzler2024krylovlocalizationprobeergodicity]. Finally, chaos is associated with delocalization in Krylov space [Dymarsky2020_QuantumChaosAsDelocalization], which can be quantified by the average position in Krylov space known as the “Krylov complexity”, the long time average of which is expected to saturate a maximum bound for chaotic systems [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Camargo2024_SpectralandKrylovComplexityBilliards].
@@ II Operator Krylov space
We will now briefly describe the Krylov subspace expansion which is explained in greater detail in Refs. [recursionmethod_vis, nandy2024quantumdynamicskrylovspace]. In the Heisenberg picture operators $O$ evolve in time under the action of a Hamiltonian $H$ as $\frac{dO}{dt}=i[H,O]$, which upon defining the Liouvillian $\mathcal{L}(\cdot)=[H,\cdot]$, allows for the formal solution $O(t)=e^{i\mathcal{L}t}(O)$. One can do a Taylor expansion of the latter which illuminates the basic idea of the recursion method: to recursively construct a basis of orthogonal operator polynomials. To define these requires defining an inner product on the operator space, and this choice is not unique. Since we’re concerned with eigenstate transitions at arbitrary energy, we choose the (normalized) Hilbert-Schmidt norm $O||^{2}=D^{-1}{\rm Tr}[O^{\dagger}O]$, where $D$ is the dimension of the many-body Hilbert space.
@@
Therefore the Lanczos coefficients $\{b_{n}\}$, which have units of energy, contain all of the information needed to compute the dynamics of the desired operator $O_{1}$. The dimension of Krylov space $D_{K}$ is related to the number of distinct frequencies and thus depends on the model being considered, however it is upper bounded at $D_{K}\leq D^{2}-D+1$ [Rabinovici2021_OperatorComplexityaJourney]. Systems with level-repulsion are expected to saturate this bound, but some special integrable systems have much smaller $D_{K}$ [Bhattacharjee2022_KComplexity_SaddleDominatedScrambling] due to their smaller dynamical Lie algebra. In absence of those dynamical symmetries the Krylov space is typically maximal, and as previously noted, the long time limit of the average position in Krylov space, otherwise known as the Krylov complexity, is often used as a measure of integrability [nandy2024quantumdynamicskrylovspace]).
In addition, as the Lanczos coefficients $b_{n}$} are calculated through repeated action of $n$ powers of $\mathcal{L}$, they are equivalent to knowing the moments of the Liouvillian $\mu_{2n}$ (explained in detail in Ref. [recursionmethod_vis]). The autocorrelation function is also simply defined in this basis as:
@@ II.1 Krylov space integrals of motion
The tri-diagonal structure of $\mathcal{L}$ in the Krylov basis gives a simple solution to the zero-mode as a function of $n$ along the semi-infinite chain in terms of the initial weight on the first site $Q_{1}$. The support on all the even $n$ vanishes, as they describe currents, while for odd $n$ one arrives at following form:
@@
In the ideal case, we’d get an exponentially localized zero-mode on the boundary of our Krylov chain, i.e. $\kappa(n)\sim n$. The latter corresponds to a perfectly dimerized chain in which all the odd couplings are greater than the even couplings. This situation essentially describes the Su–Schrieffer–Heeger (SHH) chain [SSH1979] in the topologically non-trivial insulating phase, and our integral of motion corresponds to the well known edge state. While this scenario is realized for the Majorana zero-mode in the transverse field Ising (TFI) chain [Yates2020_LifetimeofAlmostStrong, Yates2021StrongandAlmoststrongModes], it’s extremely fine tuned and unstable to small perturbations [Yeh2023SlowlyDecayingZeroMode].
There are some generic properties that highlight the difficulty of finding these types of integrals of motion. First of all, generic seed operators $O_{1}$ are expected to have matrix elements between all eigenstates of $H$ and we would therefore not expect them to have a spectral gap, e.g. in the eigenstates thermalization regime (ETH [Srednicki1994ETH]) we expect low frequency matrix elements (below the Thouless scale) to behave like random matrix theory. The SSH scenario outlined above is thus highly non-generic as it corresponds to a problem in which our initial operator $O_{1}$ has a gapped spectrum. Generically, we’d need a scenario where there is localization on the Krylov chain but the spectrum is gapless. Random, or some form of pseudo-random, Lanczos coefficients could give us that [Fleishman1977_FluctuationsandLocalizationin1D], but as we’ll discuss later this turns out to have other problems. The primary issue is that, while this results in localization, it would be hard to guarantee that the zero-mode is localized close to the boundary. In addition, for local models the Lanczos coefficients are bounded to grow at most linearly $(b_{n}\sim n)$, with a logarithmic correction in 1D $(b_{n}\sim n/\log(n))$ [Parker2019_universaloperatorgrowth, Cao_2021StatisticalmethodforOG, nandy2024quantumdynamicskrylovspace]. Typically this growth is saturated and it has the tendency to cause exponentially fast transport on the Krylov chain, further hampering the construction of integrals of motion.
@@ III Single particle localization in Krylov space
Here $c^{\dagger}_{i},c_{j}$ are fermionic creation and annihilation operators on sites $i$ and $j$ respectively, the first sum runs over nearest neighbors $\langle i,j\rangle$ with periodic boundary conditions, and the second sum runs over all sites $L^{d}$ where $d$ is the spatial dimension. The particle has a hopping amplitude $t$, which we set to $t=1$, and random on-site potentials which are constructed from i.i.d. uniformly distributed random numbers $h_{i}\in[-1,1]$. We refer to $W$ as the disorder strength. In one and two spatial dimensions this model is localized for any $W$, while in three dimensions and higher the model is initially delocalized at low disorder, but undergoes an infinite temperature metal-insulator phase transition at a critical disorder, which is estimated at $W_{c}\approx 16.5$ in three dimensions [Corrections_to_Scaling_Anderson1999, Statistics_of_Spectra_Anderson1993, SUNTAJS_2021_Spectral_properties_3D].
@@
Because we are ultimately interested in constructing the conserved quantities, which in the Anderson model are dressed local density operators, we begin our Krylov expansion with a local traceless single site density operator: $O_{1}=c^{\dagger}_{x}c_{x}-1/2$. Here the site $x$ is the site around which we’d like to construct our LIOM. We can then write the Krylov expansion of the many-particle problem completely in the single-particle operators, since the entire problem is quadratic in fermionic operators. The single particle Hamiltonian $h$ (an $L^{d}\times L^{d}$ matrix) is defined as $H=\sum_{ij}h_{ij}c^{\dagger}_{i}c_{j}$. The single-particle Green function $G$ allows us to express the Heisenberg evolved operators as $c_{x}(t)=\sum_{y}c^{\dagger}_{y}G(y,t|x,0)=\sum_{y}c^{\dagger}_{y}\left<y|e^{-iht}|x\right$>. By Wick’s theorem we arrive at following expression for the autocorrelation $C(t)$:
@@
We find that in all regimes of the three-dimensional Anderson model the fluctuations are considerably smaller than the means and decrease with $n$ and $L$, as shown in Fig. 1. In addition, the correlation matrix of the $b_{n}$ is sharply peaked only at the diagonal, so we can conceptually consider the Lanczos coefficients $b_{n}$ to be become essentially independent random variables $b_{n}$ with very small fluctuations. Consequently, typical properties of the zero-modes can be understood from the behavior of the disorder averaged Lanczos coefficients. With hindsight this behavior could have been anticipated, since with each increasing order in the Krylov expansion our operator is supported on a larger volume and therefore an increasing amount of random numbers participate in determining the operator structure.
@@
Within this model, expression (7) is not normalizable for $\gamma\leq 0$ and thus the transition occurs at $\gamma=0$. Non-decaying dimerization as in the previously discussed SSH scenario corresponds to a nonzero $\epsilon$, which also accounts for finite size corrections leading to a persistent dimerization in the extended regime, as will be shown. Thus in the thermodynamic limit or well-localized regimes we expect to find $\epsilon=0$.
We can extract the model parameters in (10) by fitting it to the coefficients shown in Fig. 2. The results for the parameters, found using a least squares fitting procedure, are shown in Fig. 3. We find $\gamma$ crosses zero around $W\approx 16$, consistent with the accepted $W_{c}\approx 16.5$ and we otherwise have $\gamma\leq 0$ in the extended regime. In addition, we find $\gamma\approx\frac{1}{2d}$ deep in the localized phase, which we predict below in Section III.3.
@@ III.1 Weak Disorder
$C_{\infty}=\lim_{t\rightarrow\infty}C(t)=|Q_{1}|^{2},$ (11)
The latter simply follows from inserting the eigenbasis of $L$ and assuming that there is only a single zero eigenvalue state. The same quantity can of course be calculated in the underlying Hamiltonian problem by taking the infinite time-average of $C(t)$, which results in

### v4

# Anderson localization: A view from Krylov space

@@ I Introduction
Modeling quenched disorder has been instrumental in understanding the electronic properties of solids. In the case of the non-interacting Anderson model, spatial disorder is known to localize the electronic states if it is larger than a critical disorder strength, whereas below this strength the electron states are extended throughout the system [3, 52, 38, 25]. Applying these concepts to interacting systems has been a major endeavor in the past decades [11, 28, 55, 53, 33, 35, 65, 72, 88, 1, 53, 2, 67, 66, 79, 37, 46, 64, 56, 9, 45, 92, 50, 47, 16, 39, 29, 22, 77, 73, 75, 16], but the interacting problem, dubbed many-body localization (MBL), has proved significantly more challenging both conceptually and numerically [85, 84, 69, 76].
@@
We may interpret these localized ergodicity-breaking systems to be effectively integrable, meaning they must possess an extensive set of local integrals of motion (LIOMs). Their existence may be motivated intuitively, invoking adiabatic evolution of the Hamiltonian from completely decoupled degrees of freedom [34], e.g. qubits, in which case LIOMs are Pauli-like operators (i.e. having discrete spectra $\pm 1$), or they can be obtained by simply time-averaging the Hamiltonian evolution of initially localized qubits [18, 27]. Though not uniquely defined, LIOMs should be exponentially localized with a small enough localization length to be stable against so called “quantum avalanches” induced by rare Griffith’s regions [17, 21, 44, 59, 51, 71, 58, 70, 19, 60, 66, 20, 86, 43, 40]. Given their non-unique construction and the stringent conditions on their decay, it is important to continue looking for new methods to understand the emergence of local integrals of motion. The Krylov subspace expansion, commonly known as the recursion method [91], is one such method which, although not a new technique, has been increasingly explored as a probe of quantum chaos (for a recent review, see Ref. [54]). The LIOMs we will explore in this work are the same operators as the time-averaged LIOMs of Ref. [18], but, importantly, expressed in the dynamically generated Krylov basis, which is arguably more natural and demonstrably more compact than any apriori fixed basis.
@@
The edge mode nature of LIOMs studied below has been identified in recent work on nearly integrable spin chains [95, 96, 97]. By way of surveying other recent efforts to exploit Krylov methods, we note there have been many studies of quantum chaos and integrability in Krylov space [57, Barbón2019OnTheEvolutionofOperatorComplexity, 61, 63, 32, 15, 6, 12, 24, 62, 30, 82, 13, 14, 49, 5, 94, 96, 97, 98, 41, 89, 7, 95, 68, 90, 87, 42, 99, 10, 8, 4] which come with various proposals for detecting ergodicity breaking transitions. For one, chaos has been associated with a maximum growth rate of the Lanczos coefficients [57], though this is not always the case [12]. The interplay between integrability, integrability breaking, and dimerization has been pointed out in Refs. [63, 30, 54, 49]. Finally, chaos is associated with delocalization in Krylov space [23], which can be quantified by the average position in Krylov space known as the “Krylov complexity”, the long time average of which is expected to saturate a maximum bound for chaotic systems [57, Barbón2019OnTheEvolutionofOperatorComplexity, 61, 63, 32, 15, 6, 12, 24, 62, 30, 13].
@@ II Operator Krylov space
We will now briefly describe the Krylov subspace expansion which is explained in greater detail in Refs. [91, 54]. In the Heisenberg picture operators $O$ evolve in time under the action of a Hamiltonian $H$ as $\frac{dO}{dt}=i[H,O]$, which upon defining the Liouvillian $\mathcal{L}(\cdot)=[H,\cdot]$, allows for the formal solution $O(t)=e^{i\mathcal{L}t}(O)$. One can do a Taylor expansion of the latter which illuminates the basic idea of the recursion method: to recursively construct a basis of orthogonal operator polynomials. To define these requires defining an inner product on the operator space, and this choice is not unique. Since we’re concerned with eigenstate transitions at arbitrary energy, we choose the (normalized) Hilbert-Schmidt norm $O||^{2}=D^{-1}{\rm Tr}[O^{\dagger}O]$, where $D$ is the dimension of the many-body Hilbert space.
@@
Therefore the Lanczos coefficients $\{b_{n}\}$, which have units of energy, contain all of the information needed to compute the dynamics of the desired operator $O_{1}$. The dimension of Krylov space $D_{K}$ is related to the number of distinct frequencies and thus depends on the model being considered, however it is upper bounded at $D_{K}\leq D^{2}-D+1$ [61]. Systems with level-repulsion are expected to saturate this bound, but some special integrable systems have much smaller $D_{K}$ [12] due to their smaller dynamical Lie algebra. In absence of those dynamical symmetries the Krylov space is typically maximal, and as previously noted, the long time limit of the average position in Krylov space, otherwise known as the Krylov complexity, is often used as a measure of integrability [54]).
In addition, as the Lanczos coefficients $b_{n}$} are calculated through repeated action of $n$ powers of $\mathcal{L}$, they are equivalent to knowing the moments of the Liouvillian $\mu_{2n}$ (explained in detail in Ref. [91]). The autocorrelation function is also simply defined in this basis as:
@@ II.1 Krylov space integrals of motion
The tri-diagonal structure of $\mathcal{L}$ in the Krylov basis gives a simple solution to the zero-mode as a function of $n$ along the semi-infinite chain in terms of the initial weight on the first site $Q_{1}$. The support on all the even $n$ vanishes, as they describe currents, while for odd $n$ one arrives at following form (illuminated further in Appendix A):
@@
In the ideal case, we’d get an exponentially localized zero-mode on the boundary of our Krylov chain, i.e. $\kappa(n)\sim n$. The latter corresponds to a perfectly dimerized chain in which all the odd couplings are greater than the even couplings. This situation essentially describes the Su–Schrieffer–Heeger (SHH) chain [81] in the topologically non-trivial insulating phase, and our integral of motion corresponds to the well known edge state. While this scenario is realized for the Majorana zero-mode in the transverse field Ising (TFI) chain [95, 96], it’s extremely fine tuned and unstable to small perturbations [97].
There are some generic properties that highlight the difficulty of finding these types of integrals of motion. First of all, generic seed operators $O_{1}$ are expected to have matrix elements between all eigenstates of $H$ and we would therefore not expect them to have a spectral gap, e.g. in the eigenstates thermalization regime (ETH [80]) we expect low frequency matrix elements (below the Thouless scale) to behave like random matrix theory. The SSH scenario outlined above is thus highly non-generic as it corresponds to a problem in which our initial operator $O_{1}$ has a gapped spectrum. Generically, we’d need a scenario where there is localization on the Krylov chain but the spectrum is gapless. Random, or some form of pseudo-random, Lanczos coefficients could give us that [26], but as we’ll discuss later this turns out to have other problems. The primary issue is that, while this results in localization, it would be hard to guarantee that the zero-mode is localized close to the boundary. In addition, for local models the Lanczos coefficients are bounded to grow at most linearly $(b_{n}\sim n)$, with a logarithmic correction in 1D $(b_{n}\sim n/\log(n))$ [57, 14, 54]. Typically this growth is saturated and it has the tendency to cause exponentially fast transport on the Krylov chain, further hampering the construction of integrals of motion.
@@ III Single particle localization in Krylov space
Here $c^{\dagger}_{i},c_{j}$ are fermionic creation and annihilation operators on sites $i$ and $j$ respectively, the first sum runs over nearest neighbors $\langle i,j\rangle$ with periodic boundary conditions, and the second sum runs over all sites $L^{d}$ where $d$ is the spatial dimension. The particle has a hopping amplitude $t$, which we set to $t=1$, and random on-site potentials which are constructed from i.i.d. uniformly distributed random numbers $h_{i}\in[-1,1]$. We refer to $W$ as the disorder strength. In one and two spatial dimensions this model is localized for any $W$, while in three dimensions and higher the model is initially delocalized at low disorder, but undergoes an infinite temperature metal-insulator phase transition at a critical disorder, which is estimated at $W_{c}\approx 16.5$ in three dimensions [78, 74, 83] and $W_{c}\approx 33.2$ in four dimensions [48]. We will focus on three dimensions in this work, as it is the smallest dimension hosting a transition for which we can explore the extended, critical, and localized regimes with the largest possible linear system size.
@@
Because we are ultimately interested in constructing the conserved quantities, which in the Anderson model are dressed local density operators, we begin our Krylov expansion with a local traceless single site density operator: $O_{1}=c^{\dagger}_{x}c_{x}-1/2$. Here the site $x$ is the site around which we’d like to construct our LIOM, ensuring that the finite amount of Krylov space operators we construct have large initial overlap with that LIOM. We can then write the Krylov expansion of the many-particle problem completely in the single-particle operators, since the entire problem is quadratic in fermionic operators. The single particle Hamiltonian $h$ (an $L^{d}\times L^{d}$ matrix) is defined as $H=\sum_{ij}h_{ij}c^{\dagger}_{i}c_{j}$. The single-particle Green function $G$ allows us to express the Heisenberg evolved operators as $c_{x}(t)=\sum_{y}c_{y}G(y,t|x,0)=\sum_{y}c_{y}\left<y|e^{-iht}|x\right$>. By Wick’s theorem we arrive at following expression for the autocorrelation $C(t)$:
@@
We find that in all regimes of the three-dimensional Anderson model the fluctuations are considerably smaller than the means and decrease with $n$ and $L$, as shown in Fig. 1 (the oscillatory behavior for $W=4$ is due to finite size and addressed in Appendix B). In addition, the correlation matrix of the $b_{n}$ is sharply peaked only at the diagonal, so we can conceptually consider the Lanczos coefficients $b_{n}$ to be become essentially independent random variables $b_{n}$ with very small fluctuations. Consequently, typical properties of the zero-modes can be understood from the behavior of the disorder averaged Lanczos coefficients. With hindsight this behavior could have been anticipated, since with each increasing order in the Krylov expansion our operator is supported on a larger volume and therefore an increasing amount of random numbers participate in determining the operator structure.
@@
The full motivation for the model will become clear throughout the remaining part of the paper. While the model doesn’t exactly describe the data, it captures all relevant features. The model is quite accurate as shown in Appendix D. In addition, the parameters are robust as shown by the small error bars in Fig. 3.
@@ III.1 Weak Disorder
$C_{\infty}=\lim_{t\rightarrow\infty}C(t)=|Q_{1}|^{2}.$ (11)
The latter simply follows from inserting the eigenbasis of $L_{n,m}$ and assuming that there is only a single zero eigenvalue state. The same quantity can of course be calculated in the underlying Hamiltonian problem by taking the infinite time-average of $C(t)$, which results in